OpenStudy (anonymous):

What happens when v is increasing (i) slightly, (ii) infinite large ? equation below:

OpenStudy (anonymous):

\(\Large{v= \sqrt{tan \theta g l sin \theta}}\)

OpenStudy (anonymous):

v=speed l=length of a string g=acceleration due to gravity 

OpenStudy (anonymous):

Your question does not make sense, but I'll try to answer it anyway. When v is increasing, the string will swing up higher, so theta increases as well.

OpenStudy (anonymous):

i know that it doesn't make sense, that's why I ask. i'm thinking that it is referring to the relationships between the speed, v, with the other variables but i'm not sure

OpenStudy (theeric):

I'm not even thinking about the physical meaning, just the relationship that's expressed by math. \(v=\sqrt{\tan(\theta)gl\cos(\theta)}\) \(g\) and \(l\) are constant. \(v=\sqrt{\tan(\theta)\cos(\theta)}\sqrt{gl}=\sqrt{\dfrac{\sin\theta}{\cancel{\cos\theta}}\cancel{\cos(\theta)}}\sqrt{gl}=\sqrt{\sin(\theta)}\sqrt{gl}\) So, \(v= \sqrt{\sin(\theta)}\sqrt{g\ l\ }\) Right? If v gets slightly bigger, \(g\) and \(l\) can get slightly bigger. Or, the \(\theta\rightarrow\dfrac\pi 2\). If it goes to \(\infty\), then you know it can't be \(\theta\)... So.... Do you see what I mean?

OpenStudy (theeric):

Let me know if you have any questions on this!

OpenStudy (anonymous):

why did you use cos theta?The equation uses sin theta.

OpenStudy (theeric):

Haha, thanks! Because I made a mistake!

OpenStudy (theeric):

@AntiNode

OpenStudy (anonymous):

but is the relationship still be the same or not? O.o

OpenStudy (theeric):

So, just don't simplify I guess! Notice that as \(\theta\rightarrow \dfrac\pi2\), \(tan\theta\rightarrow\infty\).

OpenStudy (theeric):

But the \(g\) and \(l\) can still be changing the same. No normal situation will have \(l\) or \(g\) go to \(\infty\), and I don't think that \(\theta\rightarrow\dfrac\pi2\)\(\implies v\rightarrow\infty\) in the swinging example... But, that's what the math looks like to me.

OpenStudy (anonymous):

i'm just wondering. is there a difference when the speed is increasing "SLIGHTLY" or "INFINITELY LARGE" ? because it looks same to me. Both cases are increasing so what's the difference?

OpenStudy (anonymous):

@agent0smith please?

OpenStudy (agent0smith):

Is this the actual question? Post a screenshot of it.

OpenStudy (agent0smith):

As v increases the angle increases. As v approaches infinity, the angle approaches 90 degrees.

OpenStudy (agent0smith):

@theEric that's what happens. You can try it with an object on a string - as you swing faster, the angle comes closer to being horizontal.

OpenStudy (anonymous):

only for letter c

OpenStudy (agent0smith):

Then yeah, what i said above. If v increases then sin and tan theta must also increase, since g and l are constants

OpenStudy (anonymous):

oh okay.... thanks...! one more thing... is there a difference when the speed is increasing "SLIGHTLY" or "INFINITELY LARGE" ? because the question asked it separately..

OpenStudy (anonymous):

oh nvm :) thanks!

OpenStudy (agent0smith):

I think they're just asking what happens if the speed increases a little - the angle increases infinitely large - angle approaches 90 degrees note the the first answer doesn't necessarily lead you to the second, that's why they ask both.